### v1

# HalluciBot: Is There No Such Thing as a Bad Question?

Abstract
Hallucination continues to be one of the most critical challenges in the institutional adoption journey of Large Language Models (LLMs). In this context, an overwhelming number of studies have focused on analyzing the post-generation phase - refining outputs via feedback, analyzing logit output values, or deriving clues via the outputs’ artifacts. We propose HalluciBot, a model that predicts the probability of hallucination before generation, for any query imposed to an LLM. In essence, HalluciBot does not invoke any generation during inference. To derive empirical evidence for HalluciBot, we employ a Multi-Agent Monte Carlo Simulation using a Query Perturbator to craft $n$ variations per query at train time. The construction of our Query Perturbator is motivated by our introduction of a new definition of hallucination - truthful hallucination. Our training methodology generated 2,219,022 estimates for a training corpus of 369,837 queries, spanning 13 diverse datasets and 3 question-answering scenarios. HalluciBot predicts both binary and multi-class probabilities of hallucination, enabling a means to judge the query’s quality with regards to its propensity to hallucinate. Therefore, HalluciBot paves the way to revise or cancel a query before generation and the ensuing computational waste. Moreover, it provides a lucid means to measure user accountability for hallucinatory queries.
HalluciBot: Is There No Such Thing as a Bad Question?
William Watsonthanks: Equal Contributionand Nicole Cho [^1] Equal Contribution J.P. Morgan AI Research New York, NY, USA william.watson@jpmchase.com nicole.cho@jpmorgan.com
1 Introduction
Figure 1: Comparison between traditional methods of confidence estimation and HalluciBot. In the former, user inputs a query to generate an output that is accurate or hallucinatory. If hallucinatory, user can end the session or revise the query for iterative rounds of generation. In contrast, HalluciBot predicts the probability of hallucination for a given query - before generation. Therefore, users can instantly gain insight into hallucination probability and revise a query.
Despite the promising potential for a myriad of practical use cases, Large Language Models (LLMs) offer limited insights into their chain of thought (Liang et al., 2022; Wei et al., 2023; Kojima et al., 2023; Li et al., 2023) and have the propensity to hallucinate in various circumstances (Jiang et al., 2021). Common factors that drive hallucinations encompass high model complexity, flawed data sources, or inherent sampling randomness. Specifically, the intrinsic trade-off between greedy deterministic decoding and the creativity spawned through nucleus sampling induces a heightened propensity to generate hallucinations (Huang et al., 2023). This challenge is compounded by limitations such as the frequent inaccessibility into the LLMs’ training datasets (Liang et al., 2022). HuggingFace’s release of its "Hallucinations Leaderboard" on January 29th, 2024 (Minervini et al., 2024; Gao et al., 2023) highlights the importance of resolving hallucination-related issues via the concerted effort of evaluating different LLMs. In this context, the majority of current studies have focused on the post-generation phase of output analysis as expanded in Peng et al. (2023) such as - (1) self-refinement via feedback loops on the model’s output (Madaan et al., 2023), (2) analysis of logit output values to detect hallucination (Varshney et al., 2023), or (3) for a minority of studies focused on the pre-generation phase, the ingestion of recent knowledge to improve performance (Tonmoy et al., 2024). We propose a novel model, HalluciBot, that predicts the probability of hallucination, before any generation, for any type of query. In essence, for any query posed to an LLM, HalluciBot estimates:
Figure 2: System Overview. A single query $q_{0}$, supplied by the user, is perturbed in $n$ different ways. Next, the original and perturbed queries $q_{i}$ are independently answered by the Generator agents. This Monte Carlo Simulation provides an estimate into the rate of hallucination $p_{h}(q_{0})$ for an original query $q_{0}$. Undergoing this Multi-Agent Monte Carlo Simulation, covering 13 datasets, HalluciBot is trained to predict the probability that any query $q_{0}$ will yield an hallucination, and predict the expected value of hallucinations sampled - before generation.
[…]
    a binary classification of the query’s propensity to hallucinate ("Yes" or "No"), as well as,
[…]
    a multi-class hallucination rate estimating the expected value of hallucination via sampling.
In order to understand HalluciBot, it is imperative to discuss the following key fundamental characteristics regarding hallucinations and LLMs. Firstly, hallucination itself is not a derivative of any mathematical model and intrinsically, can only be determined by juxtaposing the ground truth with the LLM’s generated outputs. Secondly, LLMs frequently advance output quality by different sampling methods such as nucleus sampling (Holtzman et al., 2020; Fan et al., 2018; Holtzman et al., 2018; Radford et al., 2018). In light of these features, HalluciBot is trained via the method of leveraging a Multi-Agent Monte Carlo Simulation for the construction of an encoder for hallucination quantification. In essence, HalluciBot, leveraging gpt-3.5-turbo, is trained via:
[…]
    perturbing 369,837 queries $n$ times into unique variations that retain the original semantic meaning yet significantly diverge lexically; then,
[…]
    employing $n+1$ independent agents to sample an output from each query (including the original) at a temperature of $1.0$ for diversity;
[…]
    conducting a Monte Carlo Simulation on 2,219,022 sampled outputs;
[…]
    deriving an empirical estimate into the expected rate of hallucination $p_{h}(q_{0})$ for the original query as the ground truth for our encoder.
As a result, our study has culminated in the following pillars of contribution.
[…]
    HalluciBot, the first encoder-based model to derive, before generation, an anticipated rate of hallucination for any type of query. HalluciBot has scored a validation accuracy of 73% on binary labels and 73% for top 3 multi-class labels.
[…]
    Moreover, our approach to construct HalluciBot absorbs the computational complexity of Monte Carlo sampling and training prior to the user session. This differs from the prevalent hallucination detection methods that involve sampling during or after the user’s current session. Thus, institutions that employ HalluciBot can systematically save on the considerable amount of computational waste engendered by "highly probable" hallucinatory queries.
[…]
    HalluciBot also generalizes to systems with Retrieval Augmented Generation (RAG) context or few-shot question answering systems with an LLM generator. Also it can generalize to systems only accessible via API calls (OpenAI, 2022; Google, 2023; Microsoft, 2023). Thus, in diverse implementation environments, HalluciBot can impart accountability to the user with regards to the query’s quality.
[…]
    We also present a myriad of behavioral and statistical discoveries from our experiments, such as an LLM’s probability to either be exceptionally correct or completely incorrect in certain scenarios. Our Appendix elaborates in detail.
2 Related Work
With regards to hallucination mitigation studies, an overwhelming majority focuses on the post-generation stage of analyzing outputs. A minority concentrates on the pre-generation phase and even amongst those, the focus lies in incorporating recent knowledge into LLMs. In detail, many expand on the universally utilized method of context-based retrieval systems (Reimers and Gurevych, 2019; Johnson et al., 2019; Nogueira and Cho, 2020; Karpukhin et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021). Other methods include relying on the model’s general knowledge (Khashabi et al., 2020) or conditioning the QA model on context generated by the LLM itself (Yu et al., 2023). Certain work has focused on mitigating hallucinations by augmenting the way LLMs generate their answers. One of the more popular techniques is to have the model enumerate its chain-of-thought (Wei et al., 2023) and think step by step (Nye et al., 2021), while building context. Another method to augment generation with context is by semantic retrieval (Lewis et al., 2021; Liu et al., 2021), handling hallucinations as they arise (Varshney et al., 2023), or using LLMs to generate context rather than retrieve (Yu et al., 2023). PromptChainer (Wu et al., 2022) profiled techniques to craft LLM chains, in which the output of one LLM’s generation process, when fed into the next LLM, can allow for more complex tasks. Language Model Cascades (Dohan et al., 2022) demonstrated that LLMs can yield probabilistic programs to tackle multi-step reasoning problems. Self-consistency (Wang et al., 2023) leveraged a new decoding strategy to sample multiple generative pathways - then select the most consistent answer. Also, Kumar et al. (2022) explored gradient-based sampling procedures that satisfy user-defined constraints. Most recent work has focused on sampling-based calibration within a single model (Cole et al., 2023) or self-verification (Kadavath et al., 2022) - the latter focuses on generating a set of outputs and feeding those back into the LLM. Furthermore, Snyder et al. (2023) explores how artifacts can differentiate hallucinated outputs. One common feature amongst these approaches is that the focus is on the output rather than the query. Also, while Zheng and Saparov (2023) study lexical perturbations, no study on hallucinations has employed a Multi-Agent approach coupled with query perturbations - which are hallmark features of HalluciBot.
3 Methodology Overview
3.1 What is Hallucination?
In general terms, hallucination refers to a false perception of patterns or objects resulting from one’s senses. With regards to LLMs, a myriad of studies have categorized hallucinations into bifurcated structures such as (1) intrinsic hallucination, which refers to the LLM’s outputs directly contradicting with the source content for extractive queries, or (2) extrinsic hallucination, which refers to the LLM’s outputs being unverifiable by the source content (i.e. irrelevant outputs) Ji et al. (2023). From a different angle, (1) factuality hallucinations refer to outputs which directly contradict or fabricate the ground truth while (2) faithfulness hallucinations define outputs that misunderstand the context or intent of the query (Huang et al., 2023).
In this study, we will focus on the aforementioned types of hallucinations and introduce truthful hallucination as a novel concept.
[…]
    The motivation for truthful hallucination stems from the analysis that neural networks display an intrinsic propensity to memorize training data (Carlini et al., 2021).
[…]
    Truthful hallucination is defined as an LLM’s inability to answer semantically similar but lexically different perturbations of a query.
[…]
    Understanding truthful hallucination is essential to account for training bias. Namely, if an LLM has memorized a query and therefore can answer it correctly, but demonstrates the inability to answer similar queries that are worded differently, it will be held accountable.
[…]
    Moreover, truthful hallucination is reflective of user behavior - unless there is a drop-down of query templates, there will almost always be diverse representations of the same query.
3.2 What is the Motivation for HalluciBot?
HalluciBot focuses on distilling LLM behavior into a speedy, encoder-based system that can predict hallucination before generation. Foremost, this is in contrast to prior work that uses multiple generations during a user’s session to provide self-consistency (Manakul et al., 2023). Next, our proposal differs from entropy based, log-prob based, or model based estimation techniques (Huang et al., 2023) that rely on the LLM’s uncertainty to predict hallucinations - these methods focus on the model’s bias while we focus on empirical estimates. Moreover, our approach consists of a Multi-Agent Simulation which stands in stark contrast to the majority of current experiments that have focused on leveraging a single LLM agent to generate outputs from a single query (Cole et al., 2023; Kadavath et al., 2022; Snyder et al., 2023). The training procedure for HalluciBot consists of:
[…]
    perturbing 369,837 queries $n$ times (§3.3)
[…]
    employing $n+1$ independent LLM agents (§3.4)
[…]
    sampling an output from each query (§3.4)
[…]
    conducting a Monte Carlo Simulation on the 2,219,022 sampled outputs (§3.5)
[…]
    training an encoder-based classifier (§3.6)
3.3 What is a Query Perturbator?
We are motivated by how perturbations:
[…]
    help comprehend truthful hallucination. We induce diversity to disentangle the generation process from any potential training bias.
[…]
    emulate real-life scenarios, in which users unintentionally create variations of the same query.
Thus, the Query Perturbator is a gpt-3.5-turbo LLM agent $\mathcal{T}$ that generates $n=5$ perturbations to the original query $q_{0}$ while retaining the same semantic meaning. Note that in practice, there are $n+1=6$ transformations because for a given query $q_{0}$, we define $T_{0}(q_{0})=I(q_{0})=q_{0}$ as an identity transform. Therefore, the original query will always be included in the set of transformed queries for the next step. In effect, the generation process can be summarized as returning a set of $\mathcal{Q}=\left\{q_{0},q_{1},\ldots,q_{n}\right\}$ query perturbations of size $n+1$. The Query Perturbator’s singular purpose is to: Rewrite the query in {$n$} radically different ways. One prompt call is sufficient to discourage duplicates. Temperature is set to $1.0$ to prioritize creativity and lexical diversity (Table 6).
$\mathcal{Q}=\mathcal{T}(q_{0})=\left[\begin{array}[]{c}I(q_{0})\\ T_{1}(q_{0})\\ \vdots\\ T_{n}(q_{0})\end{array}\right]=\left[\begin{array}[]{c}q_{0}\\ q_{1}\\ \vdots\\ q_{n}\end{array}\right]$
3.4 What is an Output Generator?
Next, the Output Generator is deployed for each perturbation due to the following reasons:
[…]
    Given the complexity of LLMs with nucleus sampling decoding, analyzing a single output from a single query may not be comprehensive enough to determine hallucination.
[…]
    Furthermore, given the risk of over-training LLMs, their opaque training data, and propensity to memorize, generating multiple outputs from the same query does not eliminate training bias. This has been a prevalent mode of studying hallucination in prior studies (§3.2)
A new generation process should be launched each time to answer the next perturbed query to ensure a truly independent output generation process. Thus, for the perturbed set $\mathcal{Q}$ for a sample $q_{0}$, the Output Generator consists of $|\mathcal{Q}|=n+1$ independent gpt-3.5-turbo LLM agents to generate outputs $a_{i}\in\mathcal{A}$ for each variation $q_{i}\in\mathcal{Q}$. The LLM agent will receive (1) for Extractive queries, a prompt with the query $q_{i}$, alongside context $c_{i}$, (2) for Multiple-Choice queries, candidate choices $k_{i}\in\mathcal{K}$, and (3) for Abstractive queries, no additional context. See Table 8 for each experiment’s prompt procedure. Temperature for all experiments is set to $1.0$ to stress-test and encourage diversity.
3.5 What is the purpose of a Monte Carlo Simulation in HalluciBot?
As evidenced by multiple studies and Section 3.1, hallucination is the outcome of multiple confounding variables - thus, it is highly unlikely that a tractable closed-form solution will be able to model hallucinations. Thus, we employ a Monte Carlo Simulation as a means to derive empirical estimations of hallucination rates in LLMs, since this method is frequently leveraged to map probability in the presence of random variable inference (Swaminathan, 2021). Thus, we estimate the probability density that a query induces hallucination.
3.6 How Do We Train a Classifier?
Once the Monte Carlo Simulation is complete for our training corpus composed of 369,837 queries spanning 13 different datasets (Appendix C, Tables 1 & 11), we start training our classifier. These queries encompass Extractive, Multiple Choice, and Abstractive scenarios. Each scenario, with or without additional context, affects the the hallucination rate of gpt-3.5-turbo. The estimates provided through our Monte Carlo Simulation are proportional to an approximation of hallucination rates.
[…]
    With a synthetic labeled set of queries $q_{0}$ and their rate of hallucinations $p_{h}(q_{0})$, we train an encoder-style classifier based on BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019) to estimate the hallucination probability density from our Monte Carlo Simulation.
[…]
    We train two versions: a binary model to estimate the propensity a query can hallucinate, and a multi-class model to predict the expected value of hallucinations if sampled $n+1$ times.
Our experiments constrain the number of perturbations to $n=5$, and when including the original query and output, we can model the hallucination rate for $n+1=6$ modes - this translates to increments of $16.6\%$ in hallucination rates.
Figure 3: Distribution of the expected value (number of hallucinations per example), rendered in percentage. For Extractive, additional context helps mitigate the rate of hallucination. For Multiple Choice, distractors can cause confusion amongst agents uniformly. However, for Abstractive, no additional information can cause massive disparities in correctness - most of our simulations resulted in no hallucinations or all hallucinations.
Figure 4: Binary distribution of labels, where at least one hallucination occurred during our simulation.
Scenario Datasets
Extractive (1) SQuADv2
Multiple TruthfulQA, SciQ, MMLU, PIQA,
Choice (9) BoolQ, OpenBookQA, MathQA
ARC - Easy, ARC - Challenge
Abstractive (6) SQuADv2, TruthfulQA, SciQ,
WikiQA, HotpotQA, TriviaQA
Table 1: Dataset Scenario Split with Reused Assets.
3.7 How Do We Encode a Query’s Scenario?
We conduct an ablation study to explore if incorporating the query’s scenario mitigates hallucinations. To create the prompt, we prepend the original query $q_{0}$ with either [EXTRACTIVE], [MULTIPLE CHOICE], or [ABSTRACTIVE], using the format <<{tag} $\mathtt{q_{0}}$}>>. Our hypothesis is based on recent research that highlights the use of RAG Yan et al. (2024); Lewis et al. (2021); Guu et al. (2020) to alleviate hallucinations. The additional context provides valuable signals related to the hallucination rate of the original query. Furthermore, we apply this technique to distinguish our experimental results from reused datasets in different scenarios, such as SciQ and SQuADv2. The queries are encoded exactly as they appear in the original datasets.
3.8 How Do We Convert Monte Carlo Estimates to Labels?
Empirical Estimate
The probability of truthful hallucinations for a query $q_{0}$, denoted as $p_{h}(q_{0})$, can be empirically estimated based on the output $a_{i}\in\mathcal{A}$ of our Multi-Agent Monte Carlo Simulation. We define the indicator function $\mathbb{I}$ to measure the incorrectness of an output $a_{i}$ with respect to the ground truth $y$ for query $q_{0}$.
$p_{h}\bigl{(}q_{0}\bigr{)}\approx\frac{1}{n+1}\sum_{i=0}^{n+1}\mathbb{I}\big{[}a_{i}\not=y\big{]}$
Binary Labels
To assess the propensity to hallucinate, we simplify the problem by considering two response values: whether $\mathit{q_{0}}$ produces any hallucination or not. Thus, we define the binary values for the probability of any hallucination as $p_{b}(q_{0})$.
$p_{b}\bigl{(}q_{0}\bigr{)}=\begin{cases}1&\text{if }p_{h}\bigl{(}q_{0}\bigr{)}>0\\ 0&\text{if }p_{h}\bigl{(}q_{0}\bigr{)}=0\\ \end{cases}$
Multi-class Labels
HalluciBot is trained to estimate the occurrence of hallucinations when queried and sampled under $n+1$ trials. To facilitate training, we convert the proportion into discrete classes by multiplying the original estimate $p_{h}(q_{0})$ by the number of agents $n+1$. This transformed variable is denoted as $\mathbb{E}{\bigl{[}p_{h}(q_{0})\bigr{]}}$.
$\mathbb{E}{\Bigl{[}p_{h}(q_{0})\Bigr{]}}=\Bigl{\lfloor}{(n+1)\cdot p_{h}(q_{0})}\Bigr{\rfloor}$
4 Metrics for Hallucination
Two key metrics are observed: accuracy and agreement. The concept of truthful hallucination focuses on the accuracy of individual agents. In addition, we analyze the level of agreement as an indicator of correlation. Moreover, the simulation’s reliability is assessed at the corpus level.
Accuracy
Accuracy serves as the measure of correctness in our evaluation by comparing the generated output $a_{i}$ to the ground truth string $y$, aiming for an almost exact match. For Multiple Choice queries, we also consider the choice label. However, due to minor variations in special characters, we adopt a partial, case-insensitive matching approach using the TheFuzz library. If there is a no match between the output $a_{i}$ and the ground truth $y$, we assign $\mathbb{I}\bigl{[}a_{i}\not=y\bigr{]}\mapsto 1$; otherwise, $\mathbb{I}\bigl{[}a_{i}=y\bigr{]}\mapsto 0$. To assess the range of correctness for each experiment, we compare the results of the original query $q_{0}$ and its output $a_{0}$ juxtaposed to the mode (most common) vote. In addition, we present the lower bound (all correct) and upper bound (at least one correct) accuracy across our $n+1=6$ agents.
Agreement
Accuracy alone is insufficient for evaluating the agreement among multiple agents when analyzing truthful hallucinations. To address this limitation, we utilize several statistical measures: Item Difficulty ($\mathbf{\mu_{D}}$) (Lord, 1952), Fleiss’s Generalized $\mathbf{\kappa}$ (Cohen, 1960; Fleiss, 1971), Mean Certainty ($\mathbf{H_{\eta}}$) (Shannon, 1948; Wilcox, 1973), and Gibbs’ $\mathbf{M_{2}}$ Index (Gibbs and Poston, 1975). These measures enable the assessment with regards to the level of agreement among independent samplings of a query, regardless of correctness. For example, if all LLM agents provide the same incorrect answer, high agreement indicates a misconception. Appendix D.3 expands in detail.
Binary Train Val Test
No $(y=0)$ 139,142 17,153 9,306
Yes $(y=1)$ 163,350 27,338 13,548
Expected Rate Train Val Test
0.0% $(y=0)$ 139,123 17,146 9,202
16.7% $(y=1)$ 35,114 4,974 2,757
33.3% $(y=2)$ 20,213 3,371 1,967
50.0% $(y=3)$ 15,749 2,757 1,768
66.7% $(y=4)$ 14,477 2,735 1,970
83.3% $(y=5)$ 17,123 3,242 2,171
100.0% $(y=6)$ 60,693 10,266 3,019
Scenario Train Val Test
Extractive 80,049 5,843 -
Multiple Choice 45,997 14,127 21,573
Abstractive 176,446 24,521 1,281
Total 302,492 44,491 22,854
Table 2: Training Splits for HalluciBot.
Reliability
We employ Cronbach’s $\alpha$ to assess the internal consistency of our Monte Carlo Simulation, as a corpus-level statistic (Cronbach, 1951).
Scenario Accuracy Agreement Rel
[…]
WAVG Extractive 85,892 92.1 91.0 69.0 97.4 87.2 85.5 84.3 75.3 99.8
Multiple Choice 81,697 76.3 76.8 47.4 91.6 71.8 75.2 71.3 61.9 92.5
Abstractive 202,248 55.9 53.9 32.9 67.3 51.2 81.5 80.0 69.1 98.8
Total 369,837 68.6 67.4 44.3 79.4 63.9 81.0 79.1 69.0 97.7
Table 3: Stage 1 Monte Carlo: Weighted average (WAVG) results are presented for each scenario and in total. The metrics are calculated using data from all examples within each scenario, regardless of the original dataset split. For a more detailed breakdown, refer to Table 15.
5 Experimental Setup
5.1 Dataset Coverage & Scenario Split
Our experiment includes 13 datasets (Table 11) divided into 3 scenarios: Extractive, Multiple Choice, and Abstractive. To evaluate the impact of context, we use SQuADv2 (Rajpurkar et al., 2016, 2018) to simulate Retrieval-Augmented Generation (Lewis et al., 2021; Guu et al., 2020). To assess the effect of multiple choice queries, we repurposed TruthfulQA (Lin et al., 2022) and SciQ (Johannes Welbl, 2017) for two experiments: one where the LLM agents select from the choices or context, and another where LLM agents generate outputs without context. Prompt templates for each output agent can be found in Table 8. We maintain the original train, validation, and test splits for each dataset across scenarios to prevent information leakage to HalluciBot.
5.2 Monte Carlo Parameters & Setup
Our Monte Carlo Simulation utilizes 2 LLM components: the Query Perturbator and Output Generator. We use gpt-3.5-turbo to sample perturbations and outputs based on the prompts provided in Table 8. All LLM agents share the same set of parameters, as described in Table 6.
Well-formedness
We used a syntactically-aware well-formedness scoring RoBERTa model (Kumar, 2021) trained on the Google Query Well-formedness Dataset (Faruqui and Das, 2018) to evaluate the grammatical correctness and completeness of 1,881,005 synthetically generated queries.
5.3 HalluciBot Parameters & Environment
We employed HuggingFace’s Trainer class with the Adam optimizer (Kingma and Ba, 2017) for training, reporting efficiency and training times in Table 9. All experiments were conducted on an AWS EC2 instance with a single GPU (Table 7).
Base Models
We fine-tune pretrained BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019) models (Table 10). To address label imbalance, we employed a weighted loss where each class weight is assigned to its inverted frequency in the training set. For the multi-class predictor, we trained a cross-entropy and ordinal model.
Train Splits
The train, validation, and test splits follow the original divisions of the datasets. Specifically, there are 302,492 training, 44,491 validation, and 22,854 testing samples. The distribution of labels across these splits is summarized in Table 2, and fine-grained splits per set are in Table 16.
[…]
Model Train Val Test Train Val Test Train Val Test Train Val Test
BERT-base-cased 80.9 64.4 66.5 81.3 68.6 72.3 86.2 74.8 74.8 76.9 63.4 70.0
+ Scenario 85.5 72.3 67.4 85.5 76.4 69.8 92.5 80.2 77.3 79.5 73.0 63.7
   + $\tau=0.129$ 81.6 71.1 67.7 84.6 79.0 75.2 77.1 71.4 69.1 93.7 88.4 82.5
RoBERTa-base 74.7 64.1 66.1 73.3 66.5 69.6 85.1 78.0 74.4 64.4 57.9 65.3
+ Scenario 79.8 73.0 69.0 79.3 76.8 71.7 88.8 81.5 78.4 71.5 72.6 66.0
   + $\tau=0.177$ 77.5 72.1 68.7 81.2 79.6 75.9 74.0 72.3 69.9 89.8 88.6 83.0
RoBERTa-large
+ Scenario 84.9 72.9 68.6 85.0 76.9 71.1 92.1 80.9 78.2 78.8 73.2 65.3
   + $\tau=0.146$ 82.5 72.5 69.2 85.0 79.5 75.3 79.0 73.4 71.8 92.0 86.8 79.3
Table 4: HalluciBot Binary Evaluation Statistics. We report the Accuracy, F1, Precision, and Recall for all data splits. Probability threshold $\tau$ is computed along the closed interval $[0,1]$ in increments of $0.001$ to maximize the F1 score. The best ablation per base model is underlined, while the overall best performing model is in bold.
[…]
Model Train Val Test Train Val Test Train Val Test Train Val Test
BERT-base-cased 49.6 32.2 24.7 69.7 49.2 40.7 81.4 62.7 56.4 70.6 53.2 49.6
+ Scenario 54.1 38.7 31.3 72.2 54.8 46.1 82.8 67.6 59.3 76.4 60.6 54.3
+ Ordinal 58.7 45.3 38.6 70.0 54.5 48.3 79.0 64.1 59.1 75.9 61.4 55.8
RoBERTa-base 47.6 34.1 26.6 66.2 50.1 42.6 77.9 62.7 57.3 67.6 54.1 51.5
+ Scenario 52.2 41.5 34.4 69.2 57.0 48.4 79.8 68.6 59.5 73.5 62.6 56.1
+ Ordinal 47.8 39.4 37.1 56.7 48.7 46.6 67.0 60.0 57.8 64.1 56.1 54.4
RoBERTa-large
+ Scenario 61.6 47.6 38.8 77.5 62.6 53.1 85.8 73.0 63.8 79.5 65.4 58.8
+ Ordinal 60.8 48.0 40.7 73.6 59.0 52.2 81.9 67.5 62.3 76.7 63.4 57.4
Table 5: HalluciBot Multi-class Evaluation Statistics. Considering the challenge of approximating a random variable and the potential presence of noise in our empirical estimate, we provide accuracy measurements for Top 1, Top 2, and Top 3 predictions. Additionally, we report the accuracy within a margin of ±$\pm 1$ of the predicted class.
6 Analysis & Discussion
6.1 Aggregated Monte Carlo Results
The aggregated Monte Carlo results demonstrate the relative performance of different question-answering scenarios. On average, the results (Table 3) indicate that Extractive outperforms Multiple Choice, which, in turn, outperforms Abstractive when subjected to perturbations. This trend suggests that the performance of gpt-3.5-turbo is influenced by the presence of additional content. Abstractive tasks show the greatest variation in agent response under perturbations, highlighting the effectiveness of added context in mitigating hallucinations (Figures 3 & 4). A full breakdown of our results is provided in Table 15.
Extractive QA
With context, gpt-3.5-turbo performs well on the SQuADv2 dataset. The mode accuracy (91.0%) and agreement (75.3%) of the agents are high. Even under radical perturbations, having unaltered context provides robustness to the agent’s capacity to answer correctly.
Multiple Choice QA
Access to answer choices mitigates hallucinations across perturbations. The ensemble accuracy is slightly higher than the baseline (+0.5%), showcasing that multiple agents can (slightly) improve accuracy rates.
Abstractive QA
When no additional context is provided, gpt-3.5-turbo achieves a mode accuracy of 53.9% under perturbations. Interestingly, there is a significant dispersion of hallucination rates compared to other scenarios (Figure 3). Moreover, there is significant variation in results among datasets. For instance, SQuADv2 shows a -59% decrease in baseline accuracy against its Extractive counterpart. In contrast, SciQ benefits in this setting, leading to a +9.4% increase in mode accuracy, as the likelihood of generating a match increases.
Coherence of Perturbations
Our analysis indicates that the perturbations created by gpt-3.5-turbo consistently exhibit a high level of coherence, as indicated by their well-formedness score of 0.87. In contrast, the original queries achieve a well-formedness score of 0.77, representing an 11.5% decline. Table 13 expands on our results, with sample perturbations in Table 14.
6.2 Can HalluciBot detect Hallucination?
Binary Results
The best HalluciBot achieves a validation accuracy of 73.0%, with a macro F1-score of 76.9% (RoBERTa-base + Scenario). With threshold tuning, improved recall and precision rates yield a F1-score of 79.6%. A full breakdown of our results is provided in Table 4.
Multi-class Results
HalluciBot achieves a validation accuracy of 47.6%, with a Top 3 accuracy of 73.0% for the RoBERTa-large + Scenario model.
Does Knowing the Scenario Matter?
Including the scenario in HalluciBot’s prompt has proven beneficial in distinguishing similar queries with varying contexts. For binary cases, incorporating the scenario leads to a notable increase of +15.5% in F1. For multi-class validation, the inclusion of scenario results in a +21.7% increase in Top 1 Accuracy and a +9.4% increase in Top 3 Accuracy.
Test Distribution Shift
The testing set poses three challenges: (1) the absence of Extractive samples; (2) an over-representation of Multiple Choice queries (94.4%); and (3) a distribution of multi-class labels that diverges with significantly fewer hallucinating-inducing queries compared to the training and validation sets. Consequently, we rely on the validation results instead.
7 Conclusion
We propose a heretofore unexplored realm of hallucination mitigation - predicting a query’s hallucination probability. HalluciBot empirically estimates how the query itself may induce hallucination. HalluciBot’s training corpus consists of diverse scenarios and domains to ensure robustness; we score an accuracy level of 73% for binary and top three multi-class labels. In essence, HalluciBot lays the foundation for a reproducible method of hallucination prediction that absorbs the cost of iterative generations during training. Moreover, the training process of HalluciBot delivers insight into the performance of gpt-3.5-turbo on lexical perturbations for 13 datasets - and present discoveries on how perturbations may induce hallucinations. Practically, institutions can implement HalluciBot to measure user accountability - penalizing those who input queries that have a hallucination probability above a certain threshold. Thus, HalluciBot’s academic and practical contributions add to the ever-growing concerted effort of enabling a robust language generation ecosystem for society.
Limitations
Despite the promising results of HalluciBot, there are limitations to be acknowledged. Firstly, our approach relies on automated LLM crowdsourcing to generate query variations, which can introduce a certain degree of variability and noise in the data. Although we took measures to ensure the quality of LLM-crowdsourced data through examining well-formedness, there may be bias or error amongst variations. Secondly, our approach is computationally expensive during training, requiring multiple LLM agents and generating $n$ variations per query. Thirdly, our datasets are English-centric, so our experimental results for robustness and consistency can only be interpreted and applied to English. Thus, it may be difficult to conclude if HalluciBot could withstand stress-testing if every query was perturbed into a new language, given that our BERT and RoBERTa backbones are English-centric. Fourthly, HalluciBot is trained soley on the outputs of gpt-3.5-turbo; however, as the framework is adaptable, future studies can train HalluciBot on any mixture of LLMs. Finally, there may be inherent bias in using gpt-3.5-turbo as both the Query Perturbator and the Output Generator. However, even though there is significant agreement regardless of correctness (Table 17), we feel confident that this is not a limiting factor of our work. Since every label is observed in significant quantities, we can assume that perturbations helped decouple semantic bias from gpt-3.5-turbo’s general knowledge.
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (“JP Morgan”), and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Appendix A Definitions
A.1 What is Extractive QA?
Extractive Question-Answering involves extracting answers directly from a given context. It can be accomplished using span-detection encoders that predict the start and end tokens of the relevant portion in the context (Devlin et al., 2018). Another approach is Retrieval-Augmented Generation (RAG), where the model generates the answer based on a selected passage containing the necessary information (Lewis et al., 2021; Guu et al., 2020).
A.2 What is Multiple Choice QA?
Multiple Choice Question-Answering is a task in few-shot learning where a model is given a question and a set of answer choices. Usually, one of the choices is correct, while the rest are distractors. In an encoder-based approach, the model replicates the question for each choice, generates a scalar response, and applies a Softmax across the choices to determine the best answer (Devlin et al., 2018). In a generative setting, the model can generate an answer choice instead of selecting from the given options, using few-shot prompting techniques (Wei et al., 2022).
A.3 What is Abstractive QA?
Abstractive Question-Answering refers to the generation of answers without access to context or candidates. Also known as Closed-book Generative QA, where no context is provided, abstractive methods are used to generate answers based solely on the question (Wei et al., 2022). Encoder models can be employed in this approach, which convert commonly occurring answer tokens into class labels (Kim et al., 2021). A multi-class model is then trained using Softmax regression to predict the answer from the question.
A.4 What is Temperature and Nucleus Sampling?
Decoder-based language models approximate the distribution of the next probable token across a given vocabulary $V$. This is often implemented by applying the Softmax function across a language model’s final output vector $u_{i}$, per token $x_{i}$.
$p\bigl{(}y_{i}\big{|}x_{1:i-1}\bigr{)}=\frac{e^{u_{i}}}{\sum_{j}e^{u_{j}}}$
Greedy decoding takes the most likely predicted token of the output distribution per sequence step, but often yields poor results and a lack of diversity. To combat this, alternative approaches leverage multinomial sampling across the distribution to sample less likely tokens. Temperature $T$ is used to control the smoothing of the distribution. As $T\to\infty$, the output distribution is smoother and more uniform; therefore unlikely tokens become more likely to sample. As $T\to 0$, the distribution approaches a Kronecker Delta ($\delta_{ij}$) function, centered on the token with the most probability mass (and mimics greedy decoding strategies). Temperature is often implemented as an adjusted Softmax function (Holtzman et al., 2020).
$p\bigl{(}y_{i}\big{|}x_{1:i-1}\bigr{)}=\frac{e^{u_{i}/T}}{\sum_{j}e^{u_{j}/T}}$
However, sampling across the full vocabulary can lead an LLM to produce extremely unlikely generations. To restrict the space of possible outcomes, two approaches exist: Top-$K$ Sampling and Nucleus Sampling. Top-$K$ sampling is implemented to include tokens that are in the Top-$K$ most probable tokens (Fan et al., 2018; Holtzman et al., 2018; Radford et al., 2018). One limitation is the static window of possible tokens considered, where unlikely tokens can still be considered even if most of the probability mass is distributed amongst fewer tokens than $K$.
Nucleus Sampling, instead of considering the Top-$K$ most probable tokens in vocabulary $V$, only considers the smallest set of tokens that have a cumulative likelihood greater than some threshold $p_{i}$. Therefore, each decoding step will consider the most likely Top-$p_{i}$ tokens $V^{(p_{i})}\subset V$, automatically eliminating improbable tokens from being accidentally sampled through masking (Holtzman et al., 2020).
$\sum_{x\in V^{(p_{i})}}p\bigl{(}x\big{|}x_{1:i-1}\bigr{)}\geq p_{i}$
A.5 What is a Multi-Agent Simulation?
In a Multi-Agent Simulation, independent agents collaborate and interact to find a solution (Russell and Norvig, 2009). In the context of our experiments, we have two types of agents: the Query Perturbator and $n+1$ Output Generators. The purpose of the Multi-Agent Simulation is to generate independent outputs after perturbing the query $q_{0}$ to observe the rate of hallucination. The presence of $n+1$ queries and agents is to ensure a balanced representation of hallucination, preventing skewed results towards either extreme. The set of outputs generated by the agents is then evaluated against the ground truth value $y$ to estimate the hallucination rate. Through multiple samplings across all 369,837 queries, we train HalluciBot to accurately identify and assess the risk of hallucination.
A.6 What is Monte Carlo Sampling?
Monte Carlo sampling is a technique used to approximate unknown quantities of interest when it’s difficult to derive an exact solution (Murphy, 2012). This is particularly useful when dealing with hidden latent variables that cannot be directly observed, such as the case of hallucination caused by interacting with a complex random variable (e.g., a Large Language Model). To address this challenge, we use Monte Carlo sampling. The idea is to conduct a Multi-Agent Simulation for each query in our training set. By doing so, we can estimate the probability of hallucination by observing the accuracy rate per query across lexical perturbations. The more simulations we conduct, the more accurate our estimation becomes. In our experiments, we sample from a series of 369,837 Multi-Agent Simulations that examine hallucination rates across perturbations for any query $q_{0}$. Therefore, we can estimate the hallucination rate $p_{h}(q_{0})$ for query $q_{0}$, given the ground truth $y$ and set of answers $a_{i}\in\mathcal{A}$.
$p_{h}\bigl{(}q_{0}\bigr{)}\approx\frac{1}{n+1}\sum_{i=0}^{n+1}\mathbb{I}\big{[}a_{i}\not=y\big{]}$
By deriving labels to estimate $p_{h}(q_{0})$, we train HalluciBot to approximate the true risk of hallucination.
A.7 What is Ordinal Regression?
Ordinal Regression, in contrast to Softmax regression, learns a sequence of cutpoints to divide the prediction space into classes, allowing models to bias errors to the nearest class label. This is useful when our labels have order, such as estimating the number of stars for a review or the expected rate of hallucination. Let $f(x)$ be an encoder-model that outputs a single scalar score, such that $f(x)\mapsto\mathbb{R}$. Additionally, let $\sigma(x)$ be the Sigmoid function:
$\sigma(x)=\frac{1}{1+\exp{\left(-x\right)}}$
Therefore, the probability of a binary ordinal classifier centered at cutoff point 00, to differentiate in log-probability space the positive and negative outcomes, is:
$p\bigl{(}y=1\big{|}x\bigr{)}$ $=\sigma\bigl{(}f(x)\bigr{)}$
$p\bigl{(}y=0\big{|}x\bigr{)}$ $=1-p\bigl{(}y=1\big{|}x\bigr{)}$
$=1-\sigma\bigl{(}f(x)\bigr{)}$
Expanding to several classes $K$, we can divide the probability space into $K-1$ cutpoints $c_{1},\dots,c_{K-1}$ with the property that $c_{1}<c_{2}<\cdots c_{K-1}$, such that the probability for each class is defined as follows (Winship and Mare, 1984):
$p\bigl{(}y=0\big{|}x\bigr{)}$ $=\sigma\bigl{(}c_{1}-f(x)\bigr{)}$
[…]
$p\bigl{(}y=k\big{|}x\bigr{)}$ $=\sigma\bigl{(}c_{k+1}-f(x)\bigr{)}$
$\quad\quad-\sigma\bigl{(}c_{k}-f(x)\bigr{)}$
[…]
$p\bigl{(}y=K-1\big{|}x\bigr{)}$ $=1-\sigma\bigl{(}c_{K-1}-f(x)\bigr{)}$
To enforce the property $c_{1}<c_{2}<\cdots c_{K-1}$ while allowing our thresholds to be differentiable, we employ a cumulative sum on a set of unbounded, learnable parameters $\theta_{1},\dots,\theta_{K-1}$, transformed by a Softplus ($\beta=1$) to avoid adding negative values; this ensures our thresholds are always increasing (Zheng et al., 2015). Therefore, each cutpoint $c_{k}$ will be of the from:
$\text{Softplus}(\theta_{k})$ $=\frac{1}{\beta}\cdot\log\bigl{(}1+\exp\left(\beta\cdot\theta_{k}\right)\bigr{)}>0$
$c_{1}$ $=\theta_{1}$
$c_{k}$ $=\theta_{1}+\sum_{i=2}^{k}\text{Softplus}(\theta_{i})$
Appendix B Additional Experiment Details
B.1 LLM Agent Settings
LLM Parameters (All Agents)
Engine gpt-35-turbo-16k
Version 2023-09-01-preview
Temperature 1.0
Frequency Penalty 0.0
Presence Penalty 0.0
Top P 0.95
Max Tokens 800
Stop None
Seed 123
Table 6: Configuration for gpt-3.5-turbo. The same set of parameters were used for both the Query Perturbator and Output Generator.
Training Instance Parameters
(All HalluciBot Models)
Instance g4dn.4xlarge
# GPUs 1
GPU Type NVIDIA T4
GPU Memory (GiB) 16
vCPUs 16
RAM (GiB) 64
Table 7: AWS EC2 training instance parameters for training HalluciBot.
Prompt Templates
Agent Prompt
Query Rewrite the query in {$n$} radically different ways.
Perturbator Query: {$q_{i}$}
Extractive You will answer the user’s query.
Output Context: {$c_{i}$}
Generator Query: {$q_{i}$}
Answer:
Multiple Choice You will answer the user’s query.
Output Query: {$q_{i}$}
Generator A) {$k_{0}$}
Z) {$k_{m}$}
Answer:
Abstractive You will answer the user’s query.
Output Query: {$q_{i}$}
Generator Short Answer:
Table 8: Prompt templates for all gpt-3.5-turbo agents. Appendix §C outlines each dataset’s taxonomy for output generation. For Multiple Choice experiments, we do not perturb any of the original choices $k_{i}\in\mathcal{K}$, and enumerate them in a consistent order for all perturbations $q_{i}\in\mathcal{Q}$. The output generator produces outputs $a_{i}\in\mathcal{A}$ for each perturbed query $q_{i}\in\mathcal{Q}$. For our experiments, the number of agents is set to $n=5$, yielding $n+1=6$ queries and outputs per example. For Extractive output generation, context is denoted as $c_{i}$.
We outline the parameters for all our LLM agents in Table 6. In addition, the prompts used per component and scenario are outlined in Table 8.
B.2 Token Usage & Statistics
We evaluated 369,837 queries, generated 1,849,185 perturbations, and generated outputs for a grand total of 2,219,022 queries. The total token usage was 717,530,842. Perturbation used 115,834,262 tokens, and the Generator used 601,696,580 tokens. Our HalluciBot is trained on 7,990,403 tokens and validated against 1,328,121 tokens, with and additional 1,305,302 tokens for testing (for BERT-base-cased). The lower token count is due to lack of prompts, context, and truncation for our HalluciBot models.
B.3 HalluciBot Training Setup & Metrics
The parameters to our AWS EC2 training instance are provided in Table 7. Additionally, we report our training metrics, including GPU hours and total floating point operations (FLOS), in Table 9. Finally, the parameters for our backbones are enumerated in Table 10.
Binary (2-Class)
GPU Time Total Update Samples Steps Train
Model Size (Hours) FLOS Steps / Second / Second Loss
BERT-base-cased 108.3M 36.3 3.98E+17 1.89E+5 11.6 1.45 0.542
+ Scenario 36.0 3.98E+17 1.89E+5 11.7 1.46 0.487
RoBERTa-base 124.6M 34.5 3.98E+17 1.89E+5 12.2 1.52 0.574
+ Scenario 36.1 3.98E+17 1.89E+5 11.6 1.46 0.518
RoBERTa-large 355.4M
+ Scenario 120.5 141E+18 1.89E+5 3.5 0.44 0.495
Multi-class (7-Class)
GPU Time Total Update Samples Steps Train
Model Size (Hours) FLOS Steps / Second / Second Loss
BERT-base-cased 108.3M 35.7 3.98E+17 1.89E+5 11.8 1.47 1.761
+ Scenario 35.7 3.98E+17 1.89E+5 11.8 1.47 1.690
+ Ordinal 36.8 3.98E+17 1.89E+5 11.4 1.43 1.778
RoBERTa-base 124.7M 34.8 3.98E+17 1.89E+5 12.1 1.51 1.798
+ Scenario 35.7 3.98E+17 1.89E+5 11.8 1.47 1.733
+ Ordinal 34.1 3.98E+17 1.89E+5 12.3 1.54 1.902
RoBERTa-large 355.4M
+ Scenario 120.2 1.41E+18 1.89E+5 3.5 0.44 1.509
+ Ordinal 117.3 1.41E+18 1.89E+5 3.6 0.45 1.622
Table 9: HalluciBot training statistics for our binary and multi-class experiments. Size is the number of learnable parameters in the model. Total FLOS is the total number of floating-point operations conducted in the model. Update Steps is the number of parameter updates the Adam optimizer (Kingma and Ba, 2017) performs on the model. We trained each model for 5 epochs in total. Training parameters can be found in Table 10. There is no difference in FLOS between BERT and RoBERTa base models as RoBERTa’s parameter increase is concentrated in a larger vocabulary.
Backbone bert-base-cased roberta-base roberta-large
Transformers Version 4.29.2 4.29.2 4.29.2
Layers 12 12 24
Attention Heads 12 12 16
Hidden Size 768 768 1024
Intermediate Size 3,072 3,072 4096
Hidden Activation GeLU GeLU GeLU
Hidden Dropout Prob 0.1 0.1 0.1
Attention Dropout Prob 0.1 0.1 0.1
Position Embedding Absolute Absolute Absolute
Precision float32 float32 float32
Max Context Length 512 514 514
Vocab Size 28,996 50,265 50,265
Total Parameters 108.3M 126.7M 355.4M
Learning Rate 5e-6 5e-6 5e-6
Warmups 0 0 0
Scheduler Type Linear Linear Linear
Weight Decay 0 0 0
Optimizer Adam Adam Adam
Adam $\beta_{1}$ 0.9 0.9 0.9
Adam $\beta_{2}$ 0.999 0.999 0.999
Adam $\epsilon_{1}$ 1e-8 1e-8 1e-8
Max Grad Norm 1.0 1.0 1.0
Training Batch Size 8 8 2
Gradient Accum. Steps 1 1 4
Number of Epochs 5 5 5
Tokenzier BertTokenizer RobertaTokenizer RobertaTokenzier
Fast True True True
Padding Strategy max_length max_length max_length
Truncation True True True
Dataset Shuffle Seed 42 42 42
Table 10: Model configuration and hyper-parameters for our HalluciBots. Configurations changed from defaults highlighted in bold. Note that for training batch size and gradient accumulation steps, we only altered roberta-large’s hyper-parameters to fit onto our GPU. All models were trained using HuggingFace’s Trainer class with the Adam optimizer (Kingma and Ba, 2017).
Appendix C Overview of Datasets Used
Domain Dataset Example Query Answer
Falsehoods TruthfulQA Which is denser, water vapor or air? Air
Physical Env. PIQA To cut a length of PVC to size, you can? Use a saw
Yes/No BoolQ Are there 51 states in the United States? No
STEM MMLU The strongest base in liquid ammonia is? $\text{NH}_{2}^{-}$
Humanities When did the first pharaohs emerge in Egypt? 5100 B.P.
Social Sciences The government measures inflation using? CPI
Science OpenBookQA What raw material is consumed by chloroplast? $\text{CO}_{2}$
SciQ Which is the final step of cell division? Cytokinesis
ARC - (C) How many valence electrons does selenium have? $6$
ARC - (E) Where is water most likely to be brackish? Estuary
Mathematics MathQA If $n=2^{0.15}$ and $n^{b}=8$ , $b$ must equal? $20$
Wikipedia SQuADv2 Where is the Mona Lisa housed? The Louvre
WikiQA What is korean money called? The won
HotpotQA EMU and Ugg boots both originated from where? Australia
General TriviaQA In an opera, whose lover was Cavaradossi? Tosca
Table 11: Overview of the 13 question-answering datasets studied in this work, the domain coverage, and examples of the question-answer format. These datasets span traditional QA formats such as Extractive, Multiple Choice, and Abstractive. Our experiments treat all scenarios as a text generation problem, albeit with different prompting templates to align responses to the ground truth answer.
Dataset License
TruthfulQA Apache-2.0
PIQA AFL-3.0
BoolQ CC BY-SA 3.0
MMLU MIT
OpenBookQA Apache-2.0
SciQ CC BY-NC 3.0
ARC - (C) CC BY-SA 4.0
ARC - (E) CC BY-SA 4.0
MathQA Apache-2.0
SQuADv2 CC BY-SA 4.0
WikiQA Other
HotpotQA CC BY-SA 4.0
TriviaQA Apache-2.0
Table 12: Licenses for each dataset. License for WikiQA is the Microsoft Research Data License Agreement for Microsoft Research WikiQA Corpus.
C.1 Extractive QA
SQuADv2 (Rajpurkar et al., 2016, 2018)
is the second version of the reading comprehension dataset from Stanford with 100,000 answerable and 50,000 unanswerable questions. Every answer is derived from a context $c_{i}$ span. For extractive QA, we only use queries that can be answered (86,821). HalluciBot is trained on 80,049 examples. For the validation set, we evaluated 5,834 items that could be answered. The reason for excluding non-answerable queries is two-fold: (1) in a zero-shot setting, it is difficult to determine if the model refuses to answer; (2) given that we are transforming the query into a semantically equivalent yet lexically distinct variation, it is impossible to determine if the new query is actually answerable from the context.
C.2 Multiple Choice QA
TruthfulQA (Lin et al., 2022)
is a QA task that gauges the "truthfulness" of LLMs. 817 questions encompass 38 categories intended to elicit false beliefs in domains such as "misconceptions" or "conspiracy theories". For the multiple choice approach, we provide all candidates with only one being correct. The Generator will then select. TruthfulQA helps measure whether our perturbed variations can act as an adversarial force to elicit false beliefs.
PIQA (Bisk et al., 2020)
is the "Physical Interaction: Question Answering" dataset focusing on physical commonsense through dichotomous choices. PIQA tests the ability of LLMs to understand how to use everyday materials; it consists of 16,113 training and 1,838 validation samples.
MMLU (Hendrycks et al., 2021b, a)
features 57 subjects from science, technology, law, humanities, and social sciences in a multiple choice format. In this experiment, each query fed to the Output Generator will be provided the perturbed query and the original answer choices. Then, it will be asked for the best choice. Accuracy is measured by the best match of the label/choice in the set of responses. There are 14,042 test, 1,531 validation, and 285 development samples.
OpenBookQA (Mihaylov et al., 2018)
tests scientific commonsense knowledge comprised of elementary science multiple choice questions from the WorldTree corpus. The train-val-test split is 4,957 training, 500 validation, and 500 testing samples.
BoolQ (Clark et al., 2019; Wang et al., 2019)
is a dichotomous QA set of yes/no questions. In our study, we exclude context and only ask for Yes/No or True/False answers. There are 9,427 training, 3,270 validation, and 3,245 test samples.
SciQ (Johannes Welbl, 2017)
contains 13,679 science questions covering Physics, Biology, and Chemistry, broken into 11,679 training, 1,000 validation, and 1,000 test samples. Each query is paired with an answer and three distractors. We omit any supporting evidence and rely solely on candidates and the Generator’s general knowledge. Since the multiple-choice answer is always consistently situated in the order of choices, we randomize the answers once before formatting. This is to enforce the model to rely on semantics instead of ordinal patterns.
ARC - Challenge (Clark et al., 2018)
is the AI2 Reasoning Challenge set consisting of 7,787 questions from grade-school level science exams. The Challenge set contains 2,590 hard questions, amalgamated from samples that were answered incorrectly by both retrieval and word co-occurrence algorithms. There are 1,119 training, 299 validation, and 1,172 test examples. Most samples have 4 answer choices, with $<1\%$ having 3 or 5 choices.
ARC - Easy (Clark et al., 2018)
is the AI2 Reasoning Easy set. It consists of 5,197 questions with a train-val-test split of 2,251, 570, and 2,376.
MathQA (Amini et al., 2019)
allows us to isolate our system on mathematical problems. The test set has no answers; therefore, our evaluations focus on the validation set of 4,475 samples.
C.3 Abstractive QA
SQuADv2 (Rajpurkar et al., 2016, 2018)
is a reading comprehension taskset (§C.1) that we repurposed for abstractive QA. By omitting the context from the Generator’s prompt, the model is conditioned on the transformed query alone. We juxtapose the differences in metrics between the extractive and abstractive settings in Table 15.
TruthfulQA (Lin et al., 2022)
as referenced in §C.2, has 817 questions to reveal misconceptions. For abstractive QA, the LLM can construct any free-text answer and match the result against the candidates via cosine similarity. The best match is either correct or a distractor.
WikiQA (Yang et al., 2015)
comprises of 3,047 questions and 29,258 sentence pairings. Although originally crafted for information retrieval tasks, we repurposed WikiQA as an abstractive task by filtering for the 1,473 sentences that were labeled as the answer to the corresponding questions. From these QA pairings, we generate answers for each perturbation. Then the labeled passage and generated sentence are aligned. If there is a high semantic similarity $(>60\%)$ between the answer and passage, the answer is considered correct.
SciQ (Johannes Welbl, 2017)
as mentioned in §C.2, evaluates 13,679 science exam questions without candidate choices. Correct responses are measured by the approximate Levenshtein distance of label substrings in the generated answer.
HotpotQA (Yang et al., 2018; Petroni et al., 2021)
consists of over 113,000 Wikipedia-based QA pairs. This set provides short, factual answers and is topically diverse. We evaluate 57,711 train samples and 5,600 validation samples. The test set contains no answers; therefore, we cannot discern any meaningful metrics.
TriviaQA (Joshi et al., 2017)
diversifies perturbations to general knowledge domains, leveraging 95,000 syntactically and lexically variable QA pairs authored by trivia enthusiasts. Only the training and validation splits contain answers; correspondingly, we evaluate 11,313 validation and 67,469 training samples.
Appendix D Metrics
We outline our metrics used in assessing the Multi-Agent Monte Carlo simulation. We also bifurcate whether metric requires labels to compute, denoted as (S)upervised and (U)nsupervised metrics.
D.1 Notation
$\begin{array}[]{rl}q_{j,0}&\text{$j$-th example, original query}\\ q_{j,i}&\text{$j$-th example, $i$-th perturbation}\\ \mathcal{Q}_{j}&\text{Set of query variations for $q_{j,i}\,\forall i$}\\ y_{j}&\text{Ground truth for query $q_{j,0}$}\\ a_{j,i}&\text{$j$-th example, $i$-th output}\\ \mathcal{A}_{j}&\text{Set of outputs for $q_{j,i}\,\forall i$}\\ n&\text{Number of perturbations}\\ n+1&\text{Number of perturbation with $q_{j,0}$}\\ m&\text{Number of examples}\\ k_{j}&\text{Number of allowed answer states}\\ f_{j,i}&\text{Frequency of $a_{j,i}$ across all $i$}\end{array}$
D.2 Accuracy (S)
Since the original query $q_{j,0}$ is in our answer set $\mathcal{A}_{j}$, we can juxtapose the baseline accuracy against ensemble metrics, such as lower-bound performance, upper-bound performance, and plurality-based accuracy. Inspired by prior work evaluating LLMs (Liang et al., 2022), we measure robustness in the worst-case (one or more raters is incorrect) and best-case (one rater is correct) scenarios. If we let $\mathbb{I}$ be the indicator function for partial matches, then let $A$ be the baseline accuracy for $n$ samples. In addition, let $\Omega$ be the lower bound, i.e. worst-case performance under perturbations $\mathcal{Q}_{j}$, and let $O$ be the upper-bound, i.e. best-case performance. Finally, given the set of $n+1$ raters, let $\hat{Y}$ be the aggregate responses by the mode of answer set $\mathcal{A}_{j}$, as a proxy for plurality voting.
$A=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}\Bigl{[}a_{j,0}=y_{j}\Bigr{]}$
$\Omega=\frac{1}{m}\sum_{j=1}^{m}\min_{i}\mathbb{I}\Bigl{[}a_{j,i}=y_{j}\Bigr{]}\leq A$
$O=\frac{1}{m}\sum_{j=1}^{m}\max_{i}\mathbb{I}\Bigl{[}a_{j,i}=y_{j}\Bigr{]}\geq A$
$\hat{Y}=\frac{1}{m}\sum_{j=1}^{m}\mathbb{I}\Bigl{[}\text{mode}\bigl{\{}\,a_{j,i}\,\bigm{|}a_{j,i}\in\mathcal{A}_{j}\,\bigr{\}}=y_{j}\Bigr{]}$
Since $q_{j,0}$ is the original query, the relationship between accuracy $A$, lower-bound performance $\Omega$, and upper-bound performance $O$ is as follows:
$\Omega\leq A\leq O$
If our raters randomly guessed, then accuracy $A$ and mode $\hat{Y}$ for $k_{j}$ choices would be $1/k_{j}$. The lower-bound $\Omega$ would approach:
$\Omega=\lim_{k_{j}\to\infty}\left(\frac{1}{k_{j}}\right)^{n+1}\approx 0$
The upper-bound performance $O$ would then be the probability of one success for $n+1$ trials.
$O=1-\left(\frac{k_{j}-1}{k_{j}}\right)^{(n+1)}$
D.3 Agreement
D.3.1 Item Difficulty (S)
For each split, the average item difficulty $\mu_{D}$ is the mean percentage of correct responses per query (Lord, 1952). It is a measure of the collective difficulty of the queries for our LLM raters. The baseline for random guessing is the expected value of a Bernoulli distribution, $\mathbb{E}\left[\mu_{D}\right]=1/k$.
$\mu_{D}=\frac{1}{m}\sum_{j=1}^{m}\Biggl{(}\frac{1}{n+1}\sum_{i=0}^{n+1}\mathbb{I}\Bigl{[}a_{j,i}=y_{j}\Bigr{]}\Biggr{)}$
D.3.2 Mean Normalized Certainty (U)
Entropy $H$ can quantify the degree of uncertainty in a set of qualitative responses. It is maximized for uniform distributions (complete uncertainty) while minimized for consistent categorizations (Shannon, 1948; Wilcox, 1973). We normalize the rater entropy $H$ by the maximum entropy allowed $H_{max}$. We reverse the scale such that 1 indicates certainty, 0 for uncertainty. Let $f_{j,i}$ denote the frequency of answer candidate $a_{i}$ for example $q_{j}$, and $k_{j}$ be the number of total allowable choices (unique states). Then proportion $p_{j,i}$ and mean normalized certainty (MNC) $H_{\eta}$ is:
$p_{j,i}$ $=\frac{f_{j,i}}{n+1}$
$H_{\eta}$ $=1-\mathbb{E}\left[\frac{H}{H_{max}}\right]$
$=1+\frac{1}{m}\sum_{j=1}^{m}\left[\sum_{i=0}^{k_{j}}\frac{p_{j,i}\log\left(p_{j,i}\right)}{\log\left(k_{j}\right)}\right]$
D.3.3 Gibbs’ M2 Index (U)
This index is a standardized metric measuring the ratio of the variance of a multinomial distribution to the variance of a binomial; since each perturbation is an independent trial, and the answer responses are categorized into exactly one of $k_{j}$ outcomes, each round of our Monte Carlo sampling is a multinomial simulation (Gibbs and Poston, 1975). Therefore, let $p_{j,i}$ be the proportion of observations for the $i$-th category and $k_{j}$ be the number of allowed categories. For readability, we reverse the index such that $M_{2}=1$ when our raters are certain and 0 when uniform.
$M_{2}=1-\frac{1}{m}\sum_{j=1}^{m}\left[\frac{k_{j}}{k_{j}-1}\left(1-\sum_{i=0}^{k_{j}}\Bigl{(}p_{j,i}\Bigr{)}^{2}\right)\right]$
D.3.4 Fleiss’s Kappa (U)
Inter-rater agreement is measured through Fleiss’ Generalized $\kappa$ (Cohen, 1960; Fleiss, 1971). This metric calculates the degree of agreement in responses over what would be expected by random chance, $1$ indicating complete agreement (and 0 for none). Let $f_{j,i}$ be the frequency of answer choice $a_{i}$ for example $q_{j}$, then the expected agreement by chance $\bar{P_{e}}$ and observed agreement $\bar{P_{o}}$ for $n+1$ raters is
$\bar{P}_{e}=\sum_{i=0}^{k_{j}}\left(\frac{1}{m(n+1)}\sum_{j=1}^{m}f_{j,i}\right)^{2}$
$P_{j}=\frac{1}{n(n+1)}\left[\left(\sum_{i=0}^{k_{j}}\Bigl{(}f_{j,i}\Bigr{)}^{2}\right)-(n+1)\right]$
$\bar{P}_{o}=\frac{1}{m}\sum_{j=1}^{m}P_{j}$
Then $\kappa$ is the ratio of the degree of agreement achieved over the degree of agreement attainable through pure chance. Note that $\kappa$ is affected by the number of raters and categories, with fewer categories often yielding higher $\kappa$ values.
$\kappa=\frac{\bar{P_{o}}-\bar{P_{e}}}{1-\bar{P_{e}}}$
D.4 Reliability: Cronbach’s Alpha (S)
For measures of internal consistency, we rely on Cronbach’s $\alpha$ (Cronbach, 1951) for dichotomous choices, in which 1 is for correct and 0 is for incorrect. We choose Cronbach’s $\alpha$ as it is widely accepted in testing theory, and is equivalent to the Kuder-Richardson Formula 20 (KR-20) for binary data (Kuder and Richardson, 1937). Let $m$ be the number of samples, $\sigma^{2}_{y}$ be each sample’s score variance across our $n+1$ raters, and $\sigma_{x}^{2}$ be the variance across the total count of correct responses per rater. Then Chronbach’s $\alpha$ is defined as:
$\alpha=\frac{m}{m-1}\left(1-\frac{\sum_{j=1}^{m}\sigma^{2}_{y}}{\sigma^{2}_{x}}\right)$
Appendix E Abusive or Sensitive Content
Throughout our experiments, when able, we captured statistics surrounding gpt-3.5-turbo’s failure to provide any response at all. While a vast majority were related to latency and servicing (APIError, ServiceUnavailableError, RateLimitError), a small subset of 2,293 samples registered AttributeError (1,081) or InvalidRequestError (1,212). The former is triggered when we generate violent or explicit content. The latter is triggered through prompt filtering, such as violent or explicit terms appearing in the prompt. In either case, we enumerate the split per dataset as follows:
$\begin{array}[]{rlrl}\text{MMLU}&638&\text{PIQA}&543\\ \text{SQuADv2}&490&\text{TriviaQA}&326\\ \text{HotpotQA}&103&\text{BoolQ}&57\\ \text{OpenBookQA}&48&\text{TruthfulQA}&29\\ \text{SciQ}&25&\text{WikiQA}&19\\ \text{MathQA}&15\\ \text{Total}&2,293\end{array}$
Appendix F Our Baseline Accuracy vs. Open-Source LLMs
Given that every original, unaltered query is answered by a gpt-3.5-turbo agent, we obtain baseline accuracy benchmarks to compare against open-source models. However, to not pollute our original discussion on our Monte Carlo Simulation and HalluciBot, we relegated the comparison here in the appendix. From our experiments, gpt-3.5-turbo outperforms all others in the HuggingFace Open LLM Leaderboard [^1] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard. For MMLU, the highest recorded accuracy is from Falcon-40b-instruct (Almazrouei et al., 2023) at 54.1% (compared to 67.6% for gpt-3.5-turbo. Also, gpt-3.5-turbo outperforms CalderaAi’s Lazarus-30B in TruthfulQA (Ours 60.4% vs. 58.3%). Finally, for the ARC Challenge and Easy set, gpt-3.5-turbo outperforms the current leader, Falcon-40b (61.9% on 25-shot) vs. our experimental results of 83.8% (C) and 92.8% (E).
When compared to benchmarks provided by LLaMA, gpt-3.5-turbo outperforms on the ARC (E) (ours 92.8% vs. 78.9%) and ARC (C) (ours 83.8% vs. 56%). However, LLaMA outperforms on PIQA (ours 80.1% vs. 82.8%) and BoolQ (ours 71.5% vs. 85.3%). For TriviaQA, gpt-3.5-turbo outperforms with 72.2% vs. 68.2% in a 0-shot environment (our experiments did not cover few-shot ablation studies, but focused on robustness under perturbations). Finally, for MMLU dataset, gpt-3.5-turbo achieved an accuracy of $67.6$ in a 0-shot setting compared with LLaMA (63.4%) in a 5-shot setting. PaLM achieved a higher 5-shot accuracy with 69.3% (Chowdhery et al., 2022).
Appendix G Sample Perturbations
We showcase 10 examples across our scenarios in Table 14.
Appendix H Well-formedness Results
We present our well-formedness results across each scenario and dataset in Table 13.
Scope Original Generated
Ex SQuADv2 0.86 0.91
Multiple Choice TruthfulQA 0.91 0.92
PIQA 0.59 0.96
MMLU 0.63 0.83
OpenBookQA 0.56 0.90
BoolQ 0.78 0.95
SciQ 0.79 0.91
ARC - Challenge 0.80 0.88
ARC - Easy 0.81 0.89
MathQA 0.33 0.75
Abstractive SQuADv2 0.86 0.91
TruthfulQA 0.91 0.93
WikiQA 0.71 0.91
SciQ 0.79 0.91
HotpotQA (KILT) 0.72 0.81
TriviaQA 0.80 0.85
Ext train 0.86 0.91
validation 0.85 0.91
MC train 0.69 0.93
validation 0.60 0.86
test 0.65 0.85
Abstr train 0.78 0.85
validation 0.80 0.86
test 0.76 0.91
Totals train 0.79 0.88
validation 0.75 0.87
test 0.65 0.85
Extractive 0.86 0.91
Multiple Choice 0.66 0.89
Abstractive 0.78 0.85
Aggregate Total 0.77 0.87
Table 13: Well-formedness scores of the original datasets juxtaposed to the average well-formedness score of gpt-3.5-turbo generated queries, with each sample averaged across the $n$ perturbed queries. We do not split our datasets between training splits as there was no significant different in scores. We color Extractive, Multiple Choice, and Abstractive scopes appropriately to differentiate scenarios, training splits, and totals.
$\mathbf{Q_{i}}$ Query
SQuADv2 $q_{0}$ How did Frederick protect Silesia when he went to invade Saxony?
$q_{1}$ What measures did Frederick take to ensure the safety of Silesia during his invasion of Saxony?
$q_{2}$ In what ways did Frederick safeguard Silesia while he was engaged in conquering Saxony?
$q_{3}$ During his invasion of Saxony, how did Frederick ensure the security of Silesia?
$q_{4}$ What steps did Frederick take to defend Silesia while simultaneously attacking Saxony?
$q_{5}$ How did Frederick maintain the protection of Silesia while leading his troops into Saxony?
PIQA $q_{0}$ How do you cut up a watermelon?
$q_{1}$ What is the proper way to slice a watermelon?
$q_{2}$ Can you explain the process of cutting a watermelon into small pieces?
$q_{3}$ Is there a certain technique to chop a watermelon effectively?
$q_{4}$ What are the steps involved in cutting a watermelon?
$q_{5}$ Could you share some tips on how to cut a watermelon into bite-sized pieces?
BoolQ $q_{0}$ Is there compulsory military service in the US?
$q_{1}$ Does the US mandate military service for its citizens?
$q_{2}$ Is military conscription mandatory in the United States?
$q_{3}$ Do US citizens have to serve in the military?
$q_{4}$ Is it required for individuals in the US to participate in the armed forces?
$q_{5}$ Is there a legal obligation for Americans to serve in the military?
TruthfulQA $q_{0}$ What did the word "posh" originally stand for?
$q_{1}$ What was the original meaning of the term "posh"?
$q_{2}$ From where did the word "posh" get its start?
$q_{3}$ Can you enlighten me about the etymology of the word "posh"?
$q_{4}$ What was the initial intention of the term "posh"?
$q_{5}$ What did "posh" signify when it first came into use?
SciQ $q_{0}$ What are found in moist forests that break down decaying plant material?
$q_{1}$ Which organisms decompose decaying plant material in damp forests?
$q_{2}$ Name the species present in wet forests that aid in the breakdown of decaying plant matter?
$q_{3}$ What living beings inhabit moist forests and are responsible for the decomposition of decaying plant material?
$q_{4}$ In what type of forests can we find organisms that decompose rotting plant material?
$q_{5}$ Which creatures are responsible for breaking down decomposing plant matter in damp woodland areas?
ARC - C $q_{0}$ Which biomolecule does not have a carbon-nitrogen bond?
$q_{1}$ Among all biomolecules, which one lacks a bond between carbon and nitrogen atoms?
$q_{2}$ Which of the biomolecules do not contain a carbon-nitrogen linkage?
$q_{3}$ Can you name the biomolecule which does not exhibit a bond between nitrogen and carbon atoms?
$q_{4}$ What is the biomolecule which doesn’t have any carbon-nitrogen bonds?
$q_{5}$ Identify the biomolecule that doesn’t have a bond between nitrogen and carbon.
MMLU $q_{0}$ A writ of certiorari from the Supreme Court indicates that the Court
$q_{1}$ The Supreme Court has issued a writ of certiorari, what does this signify?
$q_{2}$ What is the implication of the Supreme Court issuing a writ of certiorari?
$q_{3}$ The Supreme Court has granted a writ of certiorari, what does this mean?
$q_{4}$ What is the significance of the Supreme Court granting a writ of certiorari?
$q_{5}$ What does it mean when the Supreme Court issues a writ of certiorari?
WikiQA $q_{0}$ How was color introduced in film?
$q_{1}$ What is the history of incorporating color in movies?
$q_{2}$ How did the implementation of color in films come about?
$q_{3}$ What was the process behind introducing color into motion pictures?
$q_{4}$ When and how did filmmakers start using color in their productions?
$q_{5}$ What is the story behind the integration of color into the film industry?
HotpotQA $q_{0}$ What state was the man that Atchison County was named after from?
$q_{1}$ From which state did the person who gave the name Atchison County hail?
$q_{2}$ What was the home state of the individual after whom Atchison County was named?
$q_{3}$ Which state did the namesake of Atchison County belong to?
$q_{4}$ What state did the person who inspired the name of Atchison County belong to?
$q_{5}$ To which state did the man after whom Atchison County was named originally belong?
TriviaQA $q_{0}$ Which English king ruled for the shortest period?
$q_{1}$ Who is the English king with the briefest reign?
$q_{2}$ Which king of England had the shortest time in power?
$q_{3}$ Can you name the English monarch who had the quickest reign?
$q_{4}$ Which royal ruler of England had the shortest reign length?
$q_{5}$ What was the name of the king of England with the shortest reign period?
Table 14: Sample perturbations split by dataset, colored by task scenario. The Query Perturbator produces lexically distinct variations while retaining key semantic information. However, as our experiments show, variations can predispose gpt-3.5-turbo to hallucinations. Recall that $q_{0}$ is the original, unaltered query.
Datasets Accuracy Agreement Rel
[…]
Extn SQuADv2 train 80,049 91.9 90.8 68.6 97.3 87.0 85.8 84.4 75.0 99.9
val 5,843 95.2 94.1 74.5 98.8 90.6 81.2 82.7 79.3 98.3
Multiple Choice TruthfulQA val 786 60.4 60.3 39.8 76.1 58.5 88.1 79.5 72.6 37.8
PIQA train 15,677 81.2 82.3 56.7 94.0 78.8 79.1 77.6 65.1 96.8
val 1,784 80.1 83.2 58.2 93.8 79.2 79.3 77.9 66.0 82.7
MMLU dev 281 66.2 61.9 35.6 82.6 58.9 74.8 68.7 60.6 74.4
val 1,463 64.5 65.0 37.9 82.8 60.3 74.8 68.8 60.8 85.9
test 13,545 67.6 67.3 38.4 84.2 61.6 75.0 69.0 61.1 99.1
OpenBook QA train 4,909 78.0 75.6 37.8 91.3 67.8 72.8 66.6 58.0 99.1
val 497 78.1 78.9 39.8 93.0 70.5 73.4 67.5 59.2 87.1
test 499 75.6 73.7 38.1 90.8 66.8 73.1 66.9 58.4 88.9
BoolQ train 9,401 71.0 71.2 32.7 92.6 67.0 51.2 54.3 43.1 97.0
val 3,256 71.5 71.7 33.5 93.2 67.6 51.5 54.7 43.3 91.6
SciQ train 11,670 93.4 92.9 76.7 97.6 89.9 91.7 89.4 86.4 98.7
val 999 91.6 93.0 77.3 97.3 89.9 91.7 89.6 86.7 45.5
test 998 93.8 93.5 76.9 97.8 90.6 91.9 89.8 87.0 85.6
ARC - Challenge train 1,118 85.5 82.9 52.3 95.2 76.8 82.7 76.9 69.9 95.6
val 299 87.0 82.3 53.8 95.0 77.1 82.5 76.4 68.9 88.9
test 1,172 83.8 80.7 50.9 92.7 74.7 81.1 76.3 70.3 96.1
ARC - Easy train 2,248 93.3 92.7 70.8 97.9 88.0 90.1 87.4 83.7 96.6
val 570 94.4 92.1 66.1 98.1 86.7 87.6 84.6 80.8 92.4
test 2,374 92.8 92.5 69.6 98.3 87.9 90.2 86.8 82.9 95.7
MathQA train* 693 50.1 56.9 9.5 85.9 46.6 60.1 46.1 30.2 39.0
val 4,473 49.8 55.5 9.4 85.8 45.9 64.4 47.8 29.8 89.7
test 2,985 47.7 54.7 9.2 84.5 45.6 68.1 50.0 31.3 45.2
Abstractive SQuADv2 train 20,842 32.9 31.8 15.1 46.7 29.9 74.7 76.4 66.3 98.5
val 5,864 25.4 24.0 10.2 37.9 22.9 90.4 86.3 64.6 94.3
TruthfulQA val 807 52.4 28.1 61.8 78.6 55.1 58.9 61.5 53.4 31.3
WikiQA train 1,028 73.4 72.6 54.6 80.9 69.8 79.6 81.3 77.6 86.5
val 140 76.4 75.7 58.6 82.9 73.3 80.9 82.4 78.3 94.2
test 286 73.8 69.9 52.4 81.1 67.8 76.8 78.4 74.0 80.6
SciQ train 11,596 66.2 75.6 35.3 80.4 59.8 80.6 74.7 69.0 99.2
val 991 65.7 74.7 34.8 80.1 58.9 80.6 74.8 68.9 91.0
test 995 70.1 77.4 36.8 83.2 62.4 80.5 74.6 69.1 93.5
HotpotQA (KILT) train 66,345 45.5 41.7 21.2 59.6 40.6 80.7 78.6 65.8 99.8
val 5,542 42.5 38.8 21.5 55.9 38.3 72.5 74.4 69.3 96.8
TriviaQA train 76,635 71.7 69.5 48.6 79.7 66.8 84.6 83.1 72.8 99.9
dev 11,177 72.2 69.8 48.5 80.1 67.0 84.8 82.8 72.5 99.1
Table 15: Stage 1 Monte Carlo: Fine-grained experimental results for Extractive, Multiple Choice, and Abstractive question-answering scenarios. We display each dataset, split, and corresponding metrics for accuracy, ensemble accuracy (Mode), lower & upper bounds for accuracy, item difficulty ($\mathbf{\mu_{D}}$), agreement (Fleiss’s Generalized $\mathbf{\kappa}$, Mean Certainty $\mathbf{H_{\eta}}$, Gibb’s $\mathbf{M_{2}}$ Index), and reliability (Cronbach’s $\mathbf{\alpha}$). For MathQA train we only evaluated 693, out of a possible 29,800 samples.
Hallucination At least one Percent of
Datasets Rate hallucination? Hallucinated Agents per Query
Name Split No Yes No Yes 0 1 2 3 4 5 6
Extn SQuADv2 train 87.0 13.0 68.6 31.4 68.6 13.0 6.2 4.1 2.9 2.6 2.7
val 90.6 9.4 74.5 25.5 74.5 12.1 5.4 2.9 2.1 1.8 1.2
Total - 87.2 12.8 69.0 31.0 69.0 12.9 6.1 4.0 2.9 2.5 2.6
Multiple Choice TruthfulQA val 58.5 41.5 39.8 60.2 39.8 10.3 6.1 5.1 6.4 8.4 23.9
PIQA train 78.8 21.2 56.7 43.3 56.7 13.7 7.7 6.0 5.2 4.7 6.0
val 79.2 20.8 58.2 41.8 58.2 11.5 8.9 6.7 4.1 4.3 6.2
MMLU dev 58.9 41.1 35.6 64.4 35.6 13.2 5.7 6.0 11.0 11.0 17.4
val 60.3 39.7 37.9 62.1 37.9 11.3 7.5 6.6 8.2 11.3 17.2
test 61.6 38.4 38.4 61.6 38.4 11.3 8.1 7.3 8.9 10.2 15.8
OpenBook QA train 67.8 32.2 37.8 62.2 37.8 16.9 11.2 8.7 8.0 8.7 8.7
val 70.5 29.5 39.8 60.2 39.8 18.5 9.9 9.1 9.1 6.6 7.0
test 66.8 33.2 38.1 61.9 38.1 16.0 9.6 8.4 9.6 9.0 9.2
BoolQ train 67.0 33.0 32.7 67.3 32.7 18.5 13.9 10.6 9.0 7.9 7.4
val 67.6 32.4 33.5 66.5 33.5 18.7 13.2 10.7 9.0 8.0 6.8
SciQ train 89.9 10.1 76.7 23.3 76.7 9.3 4.3 2.9 2.3 2.1 2.4
val 89.9 10.1 77.3 22.7 77.3 8.4 5.0 2.6 1.7 2.3 2.7
test 90.6 9.4 76.9 23.1 76.9 11.0 4.0 1.8 2.0 2.1 2.2
ARC - Challenge train 76.8 23.2 52.3 47.7 52.3 14.4 8.9 6.8 6.4 6.4 4.8
val 77.1 22.9 53.8 46.2 53.8 11.7 10.0 6.7 8.0 4.7 5.0
test 74.7 25.3 50.9 49.1 50.9 14.2 8.0 7.2 6.1 6.2 7.3
ARC - Easy train 88.0 12.0 70.8 29.2 70.8 12.0 6.0 3.5 3.2 2.4 2.1
val 86.7 13.3 66.1 33.9 66.1 16.3 5.3 3.9 2.5 4.0 1.9
test 87.9 12.1 69.6 30.4 69.6 13.6 5.7 3.3 3.2 2.9 1.7
MathQA train 46.6 53.4 9.5 90.5 9.5 12.8 14.7 16.6 17.2 15.0 14.1
val 45.9 54.1 9.4 90.6 9.4 12.5 14.9 15.9 16.3 16.9 14.2
test 45.7 54.3 9.6 90.4 9.6 13.0 14.1 15.3 16.2 16.1 15.5
Total - 71.8 28.2 47.4 52.6 47.4 13.3 9.0 7.5 7.2 7.2 8.4
Abstractive SQuADv2 train 29.9 70.1 15.1 84.9 15.1 6.3 5.1 5.3 5.9 9.0 53.3
val 22.9 77.1 10.2 89.8 10.2 5.5 4.3 3.9 5.5 8.5 62.1
TruthfulQA val 55.1 44.9 28.1 71.9 28.1 12.9 10.9 9.9 7.2 9.5 21.4
WikiQA train 69.8 30.2 54.6 45.4 54.6 10.4 4.5 3.1 3.5 4.9 19.1
val 73.3 26.7 58.6 41.4 58.6 10.0 5.0 2.9 3.6 2.9 17.1
test 67.8 32.2 52.4 47.6 52.4 7.7 6.3 4.2 5.2 5.2 18.9
SciQ train 59.8 40.2 35.0 65.0 35.0 13.2 9.1 7.7 7.3 8.0 19.6
val 59.0 41.0 35.2 64.8 35.2 11.8 9.4 7.5 7.3 9.3 19.6
test 61.9 38.1 35.5 64.5 35.5 13.7 11.6 8.0 5.4 9.0 16.8
HotpotQA (KILT) train 40.6 59.4 21.2 78.8 21.2 9.7 6.9 6.1 6.6 9.1 40.4
val 38.3 61.7 21.5 78.5 21.5 8.0 6.0 5.3 6.4 8.9 44.1
TriviaQA train 66.8 33.2 48.6 51.4 48.6 11.7 6.0 4.5 3.9 4.9 20.3
dev 67.0 33.0 48.5 51.5 48.5 12.0 6.4 4.3 3.9 5.0 19.9
Total - 51.9 48.1 33.4 66.6 33.4 10.3 6.4 5.3 5.4 7.2 32.1
Table 16: Hallucination rates per dataset. The first two columns report the individual agent hallucination rate in total. For example, 87.0% of agents responded correctly for SQuADv2, while 13.0% hallucinated. This metric does not translate into our binary or expected values, as the latter are aggregated on a sample basis, while hallucination rates are global measures. Next, we report the binary label breakdown per split, along with the expected value labels.
[…]
Name Split # # % # % # % # %
Extn SQuADv2 train 80,049 71,560 89.4 2,037 2.5 1,121 1.4 5,331 6.7
val 5,843 5,440 93.1 120 2.1 58 1.0 225 3.9
Total - 85,892 77,000 89.6 2,157 2.5 1,1790 1.4 5,556 6.5
Multiple Choice TruthfulQA val 786 443 56.4 32 4.1 31 3.9 280 35.6
PIQA train 15,677 12,234 78.0 494 3.2 675 4.3 2,274 14.5
val 1,784 1,403 78.6 41 2.3 82 4.6 258 14.5
MMLU dev 281 164 58.4 22 7.8 10 3.6 85 30.2
val 1,463 875 59.8 68 4.6 76 5.2 444 30.3
test 13,545 8,415 62.1 746 5.5 696 5.1 3,688 27.2
OpenBook QA train 4,909 3,475 70.8 354 7.2 238 4.8 842 17.2
val 497 341 68.3 36 7.2 27 5.4 95 19.0
test 499 361 72.6 27 5.4 31 6.2 78 15.7
BoolQ train 9,401 6,238 66.4 438 4.7 458 4.9 2,267 24.1
val 3,256 2,180 67.0 147 4.5 160 4.9 769 23.6
SciQ train 11,670 10,690 91.6 205 1.8 152 1.3 623 5.3
val 999 903 90.4 12 1.2 26 2.6 58 5.8
test 998 921 92.3 15 1.5 12 1.2 50 5.0
ARC - Challenge train 1,118 899 80.4 57 5.1 28 2.5 134 12.0
val 299 241 80.6 19 6.4 5 1.7 34 11.4
test 1,172 913 77.9 69 5.9 33 2.8 157 13.4
ARC - Easy train 2,248 2,040 90.7 58 2.6 43 1.9 107 4.8
val 570 517 90.7 21 3.7 8 1.4 24 4.2
test 2,374 2,149 90.5 53 2.2 46 1.9 126 5.3
MathQA train* 693 306 44.2 41 5.9 88 12.7 258 37.2
val 4,473 1,902 42.5 325 7.3 582 13.0 1,664 37.2
test 2,985 1,252 41.9 171 5.7 382 12.8 1,180 39.5
Total - 81,697 58,862 72.0 3,451 4.2 3,889 4.8 15,495 19.0
Abstractive SQuADv2 train 20,842 5,884 28.2 980 4.7 352 1.7 13,626 65.4
val 5,864 1,236 21.1 254 4.3 79 1.3 4,295 73.2
TruthfulQA val 807 397 49.2 26 3.2 67 8.3 317 39.3
WikiQA train 1,028 720 70.0 35 3.4 15 1.5 258 25.1
val 140 104 74.3 3 2.1 2 1.4 31 22.1
test 286 194 67.8 17 5.9 6 2.1 69 24.1
SciQ train 11,596 7,035 60.7 660 5.7 371 3.2 3,530 30.4
val 991 585 59.0 62 6.3 32 3.2 312 31.5
test 995 634 63.7 58 5.8 27 2.7 276 27.7
HotpotQA (KILT) train 66,345 26,207 39.5 3,968 6.0 1,495 2.3 34,675 52.3
val 5,542 2,037 36.8 318 5.7 117 2.1 3,074 55.4
TriviaQA train 76,635 52,051 67.9 2,929 3.8 1,188 1.6 20,467 26.7
dev 11,177 7,646 68.4 427 3.8 164 1.5 2,940 26.3
Total - 202,248 104,713 51.8 9,735 4.8 3,921 1.9 83,879 41.5
- train 302,492 199,487 65.9 12,271 4.1 6,231 2.1 84,503 27.9
- val 44,491 26,270 59.0 1,905 4.3 1,525 3.4 14,791 33.2
- test 22,854 14,818 64.8 1,167 5.1 1,233 5.4 5,636 24.7
Average - - - 66.9 - 4.5 - 3.8 - 24.8
Dispersion - - - 18.9 - 1.8 - 3.2 - 17.0
Grand Total - 369,837 240,575 65.0 15,343 4.1 8,989 2.4 104,930 28.4
Table 17: We investigate the four main scenarios encountered during Stage 1 aggregation. Consensus arises when the original query and the majority of perturbed queries produce the correct answer. Dissent arises when the majority of agents disagree with the correct output generated for the original query. Corrective arises when the original query’s generated answer is incorrect, but the majority agrees on the correct answer. Finally, Erroneous arises when both the original query’s answer and the majority of agents are incorrect. The Consensus and Corrective scenarios contribute to improvements in accuracy, while the Dissent and Erroneous cases result in lower performance. As seen from the table, gpt-3.5-turbo is consistent with the majority of perturbations - regardless of correctness. This indicates that, under perturbations, gpt-3.5-turbo can make new mistakes.
# Ordinal.py
import torch
import torch.nn as nn
import torch.nn.functional as F
# Base Layer For Ordinal Prediction
class OrdinalLayer(nn.Module):
def __init__(self, n_classes, func=torch.sigmoid):
super().__init__()
self.func = func
self.theta = nn.Parameter(torch.linspace(-1, 1, n_classes - 1))
self.mask = torch.tensor([1] + [0 for _ in range(n_classes - 2)])
def forward(self, x, return_prob=False):
# B: Batch Size
# Input: x -> (B, *, 1)
size = x.size()
x = self.threshold - x.view(-1, 1)
x = torch.cat((
torch.zeros(x.size(0), 1),
self.func(x), # any cdf
torch.ones(x.size(0), 1)
), dim=-1)
x = x[:, 1:] - x[:, :-1]
# Directly gives log probs,
# Use NLL as they can not be softmaxed
# Return: Log Probs
# x -> (B, *, N_CLASSES)
if return_prob:
return x.view(*size[:-1], -1)
return (x + 1e-8).log().view(*size[:-1], -1)
@property
def threshold(self):
return (self.theta * self.mask + F.softplus(self.theta) * (1 - self.mask)).cumsum(-1)
# Wrapped Loss to avoid Softmax
class OrdinalLoss(nn.Module):
def __init__(self, **kwargs):
super().__init__()
self.loss = nn.NLLLoss(**kwargs)
def forward(self, x, y):
# x -> Logits, size: (B, C)
# y -> Labels, list (B) (like cross entropy)
return self.loss(x, y)
Listing 1: Our PyTorch implementation of an Ordinal Layer, which accepts scalar values from any model and outputs a multi-class probability distribution for n_classes.